Suppose the following bond quotes for IOU Corporation appear in the financial page of today’s newspaper. Assume the bond has a face value of $2,000 and the current date is April 19, 2018.

Company (Ticker) Coupon Maturity Last Price Last Yield EST Vol (000s)
  IOU (IOU) 5 Apr 19, 2034 106.34 ?? 1,831
a.

What is the yield to maturity of the bond? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What is the current yield? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Face Value of Bond = $2000

Coupon rate = 5% (assuming annual coupon payment)

coupon payment = $2000*5% = $100

Price of bond = $2000*106.34% = $2126.80

No of years to maturity (n) = 16 years

a). Calculating the yield of bond using IRR technique:

First, taking YTM as 4%

Price =- +.... + Coupon Payment (1 + YTM Coupon Payment Coupon Payment (1 + YTM) 2 (1+YTM) FaceValue (1 + YTM

100 100 100 2000 Price = (1 +0.04)1 + (1 +0.04)2 + .... + (1 +0.04)16 + (1 +0.04)

Price = $ 1165.23 + $ 1067.82

Price = $ 2233.05

- Secondly taking YTM as 5%

Price =- +.... + Coupon Payment (1 + YTM Coupon Payment Coupon Payment (1 + YTM) 2 (1+YTM) FaceValue (1 + YTM

100 100 100 2000 Price = (1 +0.04)1 + (1 +0.04)2 + .... + (1 +0.04)16 + (1 +0.04)

Price = $ 1083.78 + $ 916.22

Price = $ 2000

Now, Calculating YTM

PriceR- Selling Price YTM = Lower Rate+ -*(Higher Rate-Lower Rate) PriceR- PriceR

YTM = 4+ 2233.05 – 2126.8 -* (5-4) 2233,05 – 2000

YTM = 4.46%

b). Current yield = Annual Coupon Payment/Current Market Price

= 100/2126.8

= 4.70%

If you need any clarification, you can ask in comments.

If you like my answer, then please up-vote as it will be motivating